# Fast-Locking and High-Resolution Mixed-Mode DLL with Binary Search and Dead Clock Detection for Wide Frequency Ranges in 3-nm FinFET CMOS

### v1

@@ III-C Dead Clock or Toggle Detector
Although the BS algorithm significantly reduces $T_{Lock}$, it can still cause large bias overshoots due to the large step sizes at the beginning of the process, potentially leading to dead clock conditions. To address this, we implement a toggle detector circuit, shown in Fig. 7(a). This custom digital circuit consists of two flip-flops (FFs), an edge detector (ED), buffers, and an inverter. The toggle detector operates as follows: Initially, both FFs are reset. As the input clock (clkin) begins toggling, FF1 samples a constant logic ‘1’ at the rising edge of clkin. At any rising or falling edge of reflck, the ED generates a pulse that clears FF1. Since the delay between clkin and refclk is small, FF1 will be cleared before FF2 can sample logic ‘1’. This keeps the detector output (toggling) at logic ‘1’. However, if a dead clock condition occurs and refclk stops toggling, FF2 will sample the logic ‘1’, switching the toggling signal to logic ‘0’. This state persists until the FSM updates the code to a working value, at which point the detector is reset along with the PD.
@@
III-D Binary Search FSM and Controller
A simplified version of the FSM for the BS controller is illustrated in Fig. 8. During reset and initialization, the step size is set to the middle of the code range, and the initial code is zero, although both the initial step and code are configurable. The previous code (code_pre) is always stored in case a dead clock condition occurs. At each iteration, the code is either increased or decreased according to the polarity of pder, and the step size is halved by right-shifting (>>much-greater-than$>$ >) the logic ‘1’ in the step size register. Then, the toggling signal is sampled. If $toggling=1$, the BS process continues until the step size of one is reached, at which point locking is declared by raising a flag.
In the event of a dead clock condition (toggling=0), the code reverts to its last known working value, while the step size is halved to make a finer bias adjustment, and a dead_event flag is raised. During the next cycle of clkctrl the code is updated to the previous working value plus or minus the smaller step size. If the dead clock condition persists even with the finer adjustment, the FSM flags an error and enters a wait state until the DLL is reset. At this point, a new startup sequence is initiated with modified settings for the capacitor bank and the number of enabled PMOS/NMOS tails in the VCDL. A debug feature is also incorporated, enabling the FSM to freeze its operation, allowing the DAC code and clock outputs to be sampled at any given point for troubleshooting.
@@ III-D Binary Search FSM and Controller
IV Experimental Results
The proposed MM-DLL is implemented in an advanced 3-nm FinFET CMOS process, with a core area of $31.7\times 36.9~{}\mu m^{2}$ as illustrated in the optical micrograph in Fig. 10. This area includes the phase drivers that serve a whole data partition, which occupy an area of $10\times 36.9\mu m^{2}$. The MM-DLL can operate across a nominal frequency range of 533 MHz to 4.2 GHz, constrained by the internal phase-locked loop (PLL) generating clkin and clkinb rather than the DLL itself. Achieving such a low frequency in a 3-nm process requires significant capacitive loading and/or aggressive dynamic voltage scaling to increase the delay of each DE stage. The proposed MM-DLL consumes 5.4 mW at 4.2 GHz from a 0.75-V power supply. From the 5.4 mW, almost 1 mW is consumed by the phase drivers.
@@ IV Experimental Results
Figure 11: Experimental measurements of the BS locking process at 533 MHz. Transient waveforms of refclk (red) vs. fbclk (blue, with different dash patterns representing each code).
Figure 12: Experimental measurements of the BS locking process at 4.2 GHz. Transient waveforms of refclk (red) vs. fbclk (blue, with different dash patterns representing each code).
Experimental measurements of the BS locking algorithm for $f_{clkin}=533~{}MHz$ and $f_{clkin}=4.2~{}GHz$ are shown in Fig. 11 and Fig. 12, respectively. The clock signals are sampled from a digital debug chain that routes the DLL clocks to two I/O buffers, and the signals are measured on a development board using a high-speed oscilloscope configured with a 50-$\Omega$ input impedance. The phase shift between refclk (red) and fbclk (blue) progresses in a BS pattern according to the DAC code. For clarity, only five steps of the locking sequence are shown. The measured phase resolution is 0.73 ps at 4.2 GHz, which represents the minimum achievable static phase error. The measured eye diagram of fbclk is shown in Fig. 13. The measured peak-to-peak jitter ($TJ_{p2p}$) and the RMS jitter ($TJ_{RMS}$) are 14 ps and 1.2 ps, respectively. These measurements include the jitter from the digital view path (used for debugging), with no de-embedding applied. The differential input clocks (Fig. 14) show a $TJ_{p2p}$ of 9.1 ps, indicating that the proposed MM-DLL contributes only 4.9 ps of $TJ_{p2p}$. The worst-case simulated phase error between the output clock phases at 4.2 GHz is 1.5 ps.
Figure 13: Experimental measurement of fbclk (yellow) and refclk (blue), and eye diagram of fbclk.
Figure 14: Experimental measurements of input differential clocks, clkinp (yellow) and clkinn (blue), and the differential signal (gray).
@@ V Comparison
TABLE I: Performance Comparison with Previous DLLs
Reference JSSC 21’ [8] TCASI 20’ [6] TCAS 15’ [15] TVLSI 15’ [16] JSSC 16’ [12] TVLSI 21’ [20] This Work
Architecture AD-DLL MM-DLL AD-DLL AD-DLL AD-DLL AD-DLL MM-DLL
Locking scheme Coarse-fine $\Sigma\Delta$ Cyclic - locking loop TDC closed loop TDC closed loop TDC closed loop Binary - search
Technology (nm) 28 65 65 180 90 90 3
Power Supply (V) 1.0 1.2 1.1 1.8 1.2 1 0.75
Operating Frequency (GHz) 1.3-4 2.8-4 0.4-0.8 0.06-1.2 0.006–1.24 0.1–2.7 0.8–4.2
Frequency Ratio (FR) 1.02 0.35 0.67 1.81 1.98 1.86 1.55
Locking time (ns) 10 NA 51.3 27.5 9.5 53 10.5
Resolution (ps) 5.2 0.84 450 8.89 6 7.4 0.73
TJRMS / TJp2p (ps) 1.86 / 12.5 2.28 / 1.44 4.8 / 26.1 1.63 / 12.8 0.424 / 2.22 5 / 0.65 1.2 / 4.9
Efficiency (mW/GHz) 1.28 0.66 4.4 13.2 11.29 18.3 1.29
Area ($mm^{2}$) 0.0044 0.0085 0.017 0.08 0.032 0.089 0.0012
@@
FOMDLL (mW $\cdot$ V-2 $\cdot$ ps) 0.065 NA 125.8 0.55 0.23 3.87 0.011
FOMJ (mW1/2 $\cdot$ ps) 14.1 1.85 54.7 46.5 7.5 21.4 5.6
A comparison between the proposed MM-DLL and previously reported DLL architectures (with an emphasis on fast-locking DLLs) is listed in Table I. The proposed MM-DLL noticeably achieves a low locking time (10.5 ns) without relying on TDC-based locking loops. Furthermore, it delivers a high time resolution of 0.73 ps, maintaining a high power efficiency of 1.29 mW/GHz, which is competitive with or better than prior designs [8, 15, 16, 21]. The $TJ_{p2p}$ and $TJ_{RMS}$ of the proposed MM-DLL stand out compared to previous works. The jitter is expected to further improve if the MM-DLL is experimentally evaluated as a standalone component and not within an SoC as in this work.
To compare the operating range of the different DLLs, a frequency-range (FR) ratio is defined as in [22]:
$FR=\frac{2(f_{max}-f_{min})}{f_{max}+f_{min}},$ (6)
where $f_{max}$ and $f_{min}$ represent the maximum and minimum operating frequencies, respectively. Three figure-of-merits (FOMs) are defined for fair comparison between the different works. Firstly, a power FOM is defined as in [20]:
$FOM_{P}=\frac{P}{V_{CC}^{2}\cdot f_{max}\cdot FR}~{}[mW\cdot V^{-2}\cdot GHz^{-1}],$ (7)
where $P$ is the power consumption, and $V_{CC}$ is the power supply voltage. Secondly, a FOM which takes into account the locking time and resolution is defined as:
$FOM_{DLL}=FOM_{P}\cdot T_{Lock}\cdot\Delta T_{LSB}~{}[mW\cdot V^{-2}\cdot ps].$ (8)
Finally, a FOM taking into account jitter is defined similar to [22]:
$FOM_{J}=TJ_{RMS}\cdot\sqrt{\frac{P}{f_{max}}}~{}~{}[mW^{1/2}\cdot GHz^{-1/2}\cdot ps].$ (9)
The proposed MM-DLL achieves a FOMP of $1.48~{}mW\cdot V^{-2}\cdot GHz^{-1}$, which is comparable to or better to that of the prior art [12, 9, 8, 21, 23, 6, 20]. This is a major achievement, considering that they are AD-DLLs. Furthermore, thanks to its high time resolution and BS locking algorithm, the proposed MM-DLL architecture achieves a FOMDLL of $0.011~{}mW\cdot V^{-2}\cdot ps$, better than previous proposed architectures. The $T_{Lock}$ vs. resolution for various DLL architectures is shown in Fig. 15. The proposed MM-DLL outperforms previous designs by achieving both high resolution and fast locking time, effectively breaking the conventional trade-off between these two critical parameters. The FOMJ obtained is $5.6~{}mW^{1/2}\cdot GHz^{-1/2}\cdot ps$, second only to [6]. Note that the $TJ_{RMS}$ and $TJ_{p2p}$ measurements in the proposed MM-DLL include the jitter of the digital view path.
Figure 15: Locking time vs. resolution tradeoff for different DLL architectures.
@@ VI Conclusion
This paper presents a wide-range, fast-locking MM-DLL implemented in a 3-nm FinFET CMOS process. While prior designs have often traded off timing resolution for faster locking, the proposed DLL leverages a BS controller to reduce locking time without compromising resolution or jitter. This architecture achieves a startup locking time of just eleven control clock cycles, improving overall system latency. To address potential dead clock conditions caused by bias overshoot, a toggle detector circuit is integrated into the control loop, enabling the BS FSM to recover the last functional DAC code and adjust the step size. The high-resolution DAC and BBPD further enable sub-ps time resolution and low jitter. With its combination of fast locking, high resolution, and robust dead clock recovery, the proposed MM-DLL is well suited for high-performance parallel I/O interfaces and other clock-sensitive applications in modern systems-on-chip.

### v2

@@ III-D Clock Failure Detector
Although the BS algorithm significantly reduces $T_{Lock}$, it can still cause large bias overshoots due to the large step sizes at the beginning of the process, potentially leading to clock failure conditions. To address this, we implement a toggle detector circuit, shown in Fig. 7(a). This custom digital circuit consists of two flip-flops (FFs), an edge detector (ED), buffers, and an inverter. The toggle detector operates as follows: Initially, both FFs are reset. As the input clock (CLKIN) begins toggling, FF1 samples a constant logic ‘1’ at the rising edge of CLKIN. At any rising or falling edge of CLKREF, the ED generates a pulse that clears FF1. Since the delay between CLKIN and CLKREF is small, FF1 will be cleared before FF2 can sample logic ‘1’. This keeps the detector output (toggling) at logic ‘1’. However, if a stalled clock condition occurs and CLKREF stops toggling, FF2 will sample the logic ‘1’, switching the toggling signal to logic ‘0’. This state persists until the FSM updates the code to a working value, at which point the toggle detector is reset along with the BBPD.
@@
III-E Binary Search FSM and Controller
A simplified version of the FSM for the BS controller is illustrated in Fig. 8. During reset and initialization, the step size is set to the middle of the code range, and the initial code is zero, although both the initial step and code are configurable. The previous code (codepre) is always stored in case a failure clock condition occurs. At each iteration, the code is increased or decreased according to the polarity of PDER, and the step size is halved by right-shifting (>>much-greater-than$>$ >) the logic ‘1’ in the step size register. Then, the toggling signal is sampled. If $toggling=1$, the BS process continues until the step size of one is reached, at which point locking is declared by raising a flag.
In the event of a stalled clock condition (toggling=0), the code reverts to its last known working value, while the step size is halved to make a finer bias adjustment, and a stall_event flag is raised. During the next cycle of CLKCTRL the code is updated to the previous working value plus or minus the smaller step size. If the stalled clock condition persists even with the finer adjustment, the FSM flags an error and enters a wait state until the DLL is reset. At this point, a new startup sequence is initiated with modified settings for the capacitor bank and the number of enabled PMOS/NMOS tails in the VCDL. A debug feature is also incorporated, enabling the FSM to freeze its operation, allowing the DAC code and clock outputs to be sampled at any given point for troubleshooting.
@@ III-E Binary Search FSM and Controller
IV Performance Evaluation and Experimental Results
The MM-DLL is implemented in an advanced 3-nm FinFET CMOS process, with a core area of $31.7\times 36.9~{}\mu m^{2}$ as illustrated in the optical micrograph in Fig. 10. This area includes the phase drivers that serve a whole data partition, which occupy an area of $10\times 36.9\mu m^{2}$. The MM-DLL can operate across a nominal frequency range of 533 MHz to 4.26 GHz, constrained by the internal phase-locked loop (PLL) generating CLKIN,P and CLKIN,N rather than the DLL itself. Achieving such a low frequency in a 3-nm process requires significant capacitive loading and/or aggressive dynamic voltage scaling to increase the delay of each DE stage. The MM-DLL consumes 5.4 mW from a 0.75-V power supply at $f_{clkin}=4.26~{}GHz$. A power breakdown is presented in Fig. 11. As expected, the VCDL accounts for almost 80% of the power consumption. From the 5.4 mW, almost 1 mW is consumed by the phase drivers within the VCDL.
@@ IV Performance Evaluation and Experimental Results
Figure 11: Power breakdown of the proposed MM-DLL.
Figure 12: Experimental measurements of the BS locking process at 533 MHz. Transient waveforms of CLKREF (red) vs. CLKFB (blue, with different dash patterns representing each code). Locking is achieved at code 957.
Figure 13: Experimental measurements of the BS locking process at 4.26 GHz. Transient waveforms of CLKREF (red) vs. CLKFB (blue, with different dash patterns representing each code). Locking is achieved at code 725.
Figure 14: Simulation results of the eight output clock phases (CLKOUT[7:0]) for $f_{clkin}=4.26~{}GHz$. The maximum phase error betweem output clocks is 0.8 ps.
Experimental measurements of the BS locking algorithm for $f_{clkin}=533~{}MHz$ and $f_{clkin}=4.26~{}GHz$ are shown in Fig. 12 and Fig. 13, respectively. The clock signals are sampled from a digital debug chain that routes the DLL clocks to two I/O buffers, and the signals are measured on a development board using a high-speed oscilloscope configured with a 50-$\Omega$ input impedance. The phase shift between CLKREF (red) and CLKFB (blue) progresses in a BS pattern according to the DAC code. For clarity, only five steps of the locking sequence are shown. The measured phase resolution is 0.73 ps at 4.26 GHz, which represents the minimum achievable static phase error. The measured eye diagram of CLKFB is shown in Fig. 15. The measured peak-to-peak jitter ($TJ_{p2p}$) and the RMS jitter ($TJ_{RMS}$) are 14 ps and 1.2 ps, respectively. These measurements include the jitter from the digital view path (used for debugging), with no de-embedding applied. The differential input clocks (Fig. 16) show a $TJ_{p2p}$ of 9.1 ps, indicating that the proposed MM-DLL contributes only 4.9 ps of $TJ_{p2p}$. The simulated phase error between the output clock phases at 4.26 GHz is 0.8 ps, as shown in Fig. 14.
Figure 15: Experimental measurement of CLKFB (yellow) and CLKREF (blue), and eye diagram of CLKFB. Measured $TJ_{p2p}$ and $TJ_{RMS}$ are 14 ps and 1.23 ps, respectively.
Figure 16: Experimental measurements of input differential clocks, CLKIN,P (yellow) and CLKIN,N (blue), and the differential signal (gray). The measured $TJ_{p2p}$ and $TJ_{RMS}$ are 9.1 ps and 965.5 fs, respectively.
@@ V Comparison
TABLE I: Performance Comparison with Prior DLL Architectures
Reference TCASI 20’ [7] JSSC 21’ [9] TCAS 15’ [16] TVLSI 15’ [17] JSSC 16’ [13] TVLSI 21’ [22] This Work
Architecture MM-DLL AD-DLL AD-DLL AD-DLL AD-DLL AD-DLL MM-DLL
Locking scheme $\Sigma\Delta$ Coarse-fine Cyclic - locking loop TDC closed loop TDC closed loop TDC closed loop Binary - search
Technology (nm) 65 28 65 180 90 90 3
Power Supply (V) 1.2 1.0 1.1 1.8 1.2 1 0.75
Operating Frequency (GHz) 2.8-4 1.3-4 0.4-0.8 0.06-1.2 0.006–1.24 0.1–2.7 0.533–4.26
Frequency Ratio (FR) 0.25 1.02 0.67 1.81 1.98 1.86 1.55
Locking time (ns) 1300 10 51.3 27.5 9.5 53 10.5
Resolution (ps) 0.82 5.2 450 8.89 6 7.4 0.73
TJRMS / TJp2p (ps) 0.84 / 2.28 1.86 / 12.5 4.8 / 26.1 1.63 / 12.8 0.424 / 2.22 0.65 / 5 1.2 / 4.9
Efficiency (mW/GHz) 0.74 1.28 4.4 13.2 11.29 18.3 1.27
Area (mm2) 0.0085 0.0044 0.017 0.08 0.032 0.089 0.0012
FOMP (pJ) 2.86 1.25 6.6 7.25 5.89 9.85 0.82
FOMJ (pJ $\cdot$ ps) 1.6 15.9 114.8 168.9 25.1 91.5 6.2
FOMLR (pJ $\cdot$ ns2) 0.76 0.066 101.5 3.2 0.65 7.2 0.010
A comparison between the proposed MM-DLL and previously reported DLL architectures (with an emphasis on fast-locking DLLs) is listed in Table I. The proposed MM-DLL in 3-nm FinFET CMOS noticeably achieves a low locking time of 10.5 ns, without relying on TDC-based locking loops. Furthermore, it delivers a state-of-the-art time resolution of 0.73 ps, maintaining a high power efficiency of 1.27 mW/GHz, second only to [7]. Notably, the phase drivers contribute with $\sim$1 mW to the total power, and the parasitic capacitance from the capacitor bank, necessary for low-frequency operation, slightly degrades power efficiency. The proposed MM-DLL also demonstrates outstanding $TJ_{p2p}$ and $TJ_{RMS}$ performance compared to previous works. These figures are expected to improve further if the MM-DLL is characterized as a standalone component rather than within an SoC, as the current jitter measurements include contributions from the digital view path.
A frequency-range (FR) ratio is defined as in [23]:
$FR=\frac{2(f_{max}-f_{min})}{f_{max}+f_{min}},$ (8)
where $f_{max}$ and $f_{min}$ represent the maximum and minimum operating frequencies, respectively. The proposed MM-DLL achieves a large FR of 1.55. Three figure-of-merits (FOMs) are defined for a fair comparison between the different works. Firstly, a power FOM is defined similar to [22]:
$FOM_{P}=\frac{P}{f_{max}\cdot FR}~{}[J],$ (9)
where $P$ is the power consumption. Secondly, a FOM which considers the locking time and resolution is defined as:
$FOM_{LR}=FOM_{P}\cdot T_{Lock}\cdot\Delta T_{LSB}~{}[J\cdot s^{2}].$ (10)
Finally, a FOM taking into account jitter is defined as:
$FOM_{J}=TJ_{p2p}\cdot\frac{P}{f_{max}}~{}~{}[J\cdot s].$ (11)
@@
Figure 17: Locking time vs. resolution tradeoff for different DLL architectures. The proposed MM-DLL breaks the conventional trade-off between $T_{Lock}$ and resolution.
@@ VI Conclusion
This paper presents a wide-range, fast-locking MM-DLL implemented in a 3-nm FinFET CMOS process. Although prior designs have often traded off timing resolution for faster locking, the proposed DLL leverages a BS controller to reduce locking time without compromising resolution or jitter. This architecture achieves a startup locking time of just eleven control clock cycles, improving overall system latency. To address potential clock failure conditions caused by bias overshoot, a toggle detector circuit is integrated into the control loop, enabling the BS FSM to recover the last functional DAC code and adjust the step size. The high-resolution DAC and BBPD further enable sub-ps time resolution and low jitter. With its combination of fast locking, high resolution, and robust clock failure recovery, the proposed MM-DLL is well suited for high-performance parallel I/O interfaces, such as DDR and D2D links, and other clock-sensitive applications in modern SoCs.